You purchased 100 shares of stock for $10 dollars per share. After holding the stock for 10 years and not receiving any dividends you sell the stock for $42 per share. what is the effective annual rate of return on this investment?
Answer #1

Ans 15.43%

Effective annual rate of return on this investment = 15.43%

FV = Future Value PV = Present Value r= rate of interest n= no of period FV/ PV = (1+r)^n 42 / 10 = (1 +r%) 10 4.20 = (1 +r%)